OpenStudy (anonymous):

Express √-72 in i notation. A.6i√-2 B.72i C.6i√2 D.√-72i

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

\(\sqrt{-72} = \sqrt{-1} \times \sqrt{72} = i \sqrt{72}\)

ganeshie8 (ganeshie8):

you can simplify further